OpenStudy (anonymous):

Define Euler’s Phi function. Find ϕ(360), ϕ(1001) and ϕ(5040).

OpenStudy (anonymous):

I will do 360 for you. First 360=2^3 * 3^2 *5 So phi(360)=(2-1)*2^(3-1) * (3-1)*3^(2-1)*(5-1)*5^(1-1)=1*2^2 *2*3^1*4*5^0=96

OpenStudy (amoodarya):